### v3

# Eyes Can Deceive: Benchmarking Counterfactual Reasoning Capabilities of Multimodal Large Language Models

Abstract
Counterfactual reasoning, as a crucial manifestation of human intelligence, refers to making presuppositions based on established facts and extrapolating potential outcomes. Existing multimodal large language models (MLLMs) have exhibited impressive cognitive and reasoning capabilities, which have been examined across a wide range of Visual Question Answering (VQA) benchmarks. Nevertheless, how will existing MLLMs perform when faced with counterfactual questions? To answer this question, we first curate a novel CounterFactual MultiModal reasoning benchmark, abbreviated as CFMM, to systematically assess the counterfactual reasoning capabilities of MLLMs. Our CFMM comprises six challenging tasks, each including hundreds of carefully human-labeled and GPT-generated counterfactual questions, to evaluate MLLM’s counterfactual reasoning capabilities across diverse aspects. Through experiments, interestingly, we find that existing MLLMs prefer to believe what they see, but ignore the counterfactual presuppositions presented in the question, thereby leading to inaccurate responses. Furthermore, we evaluate a wide range of prevalent MLLMs on our proposed CFMM. The significant gap between their performance on our CFMM and that on several VQA benchmarks indicates that there is still considerable room for improvement in existing MLLMs toward approaching human-level intelligence. On the other hand, through boosting MLLMs performances on our CFMM in the future, potential avenues toward developing MLLMs with advanced intelligence can be explored.
1 Introduction
Figure 1: The comparison of our proposed CFMM with existing VQA benchmarks in terms of dataset scale and the degree of approaching the human-level intelligence. We categorize all VQA benchmarks into three groups based on their proximity to human-level intelligence and distinguish between these groups using different markers. We also provide examples within each group for intuitive understanding.
Recently, the Multimodal Large Language Models (MLLMs) (Achiam et al. 2023; Alayrac et al. 2022; Bai et al. 2023; Dai et al. 2024; Li et al. 2023d, e; Peng et al. 2023; Jiao et al. 2024; Zhu et al. 2023) have been a rising research hotspot due to their remarkable content comprehension and logic reasoning capabilities. Harnessing the Large Language Model (LLM) as the brain, MLLM bridges the visual world and language space, empowering the model to observe the vast world and enrich its mastered knowledge. Pioneering works like Qwen-VL (Bai et al. 2023), LLaVA (Liu et al. 2024) and MiniGPT-4 (Zhu et al. 2023) pursue GPT-4V-like (Achiam et al. 2023) visual perception and comprehension abilities by performing multimodal instruction tuning. Subsequent approaches, including Kosmos-2 (Peng et al. 2023), Shikra (Chen et al. 2023), Lumen (Jiao et al. 2024), etc., further enhance the visual comprehension by endowing the model with fine-grained grounding capabilities.
As diverse MLLMs continue to rapidly evolve, benchmarks evaluating the versatile emergent capabilities of MLLMs are concurrently being developed to keep pace with their advancements. Early benchmarks like VQAv2 (Goyal et al. 2017), COCO-Caption (Lin et al. 2014), GQA (Hudson and Manning 2019), OK-VQA (Marino et al. 2019), etc., focus on assessing visual content comprehension capabilities on specific domains. In addition, the evaluation metrics they employed can not comprehensively reflect model performances. To sufficiently assess the emergent capabilities of existing MLLMs, a new generation of benchmarks (Fu et al. 2023a; Li et al. 2023a; Liu et al. 2023) has been introduced. These benchmarks encompass a broader range of questions and employ more reliable evaluation metrics. A key advancement of these benchmarks is their emphasis on requiring the synergy of vision and language modalities for more complex reasoning to derive answers. However, the reasoning capabilities they addressed still lag far behind real human intelligence.
One of the most pivotal aspects of human intelligence lies in its capacity for imagination. Building upon the perceived world and mastered knowledge, imagination empowers humans to explore hypothetical scenarios and potential outcomes beyond the confines of reality. This cognitive process is referred to as the “counterfactual reasoning” in psychology (Epstude and Roese 2008; Van Hoeck, Watson, and Barbey 2015). Essentially, a counterfactual question can be formulated as: “Given F, if P happens, how will Q unfold?”. Here, F represents the known facts, P indicates the imaginary presupposition of changing the status of F, and Q refers to queried concepts like numbers, status, actions, etc. In the scenario of multimodal content comprehension, images are generally employed as F to reflect the real world, while P and Q are expressed with natural language. Interestingly, we observe that Although MLLMs demonstrate strong performance on prevalent VQA benchmarks, they still struggle to tackle these counterfactual questions. Through further analysis, we find that existing MLLMs are prone to believe what they see, but are lazy to comprehend the presupposition presented in sentences, thus leading to incorrect responses. These results indicate that there is still significant room for improvement for MLLMs to approach the level of human intelligence.
To assess the advanced counterfactual reasoning capabilities of existing MLLMs and explore potential avenues for further optimizing MLLMs, we introduce CFMM, a CounterFactual MultiModal reasoning benchmark in this paper. Our CFMM comprises a suite of 6 tasks to comprehensively evaluate model’s counterfactual reasoning capabilities in diverse aspects. As shown in Fig 1, compared with prevalent VQA benchmarks, our CFMM presents higher demands on the MLLM’s intelligence and encompasses substantial amounts of data samples. As for the dataset curation, firstly, to avoid the data leakage problem, we collect 1,200 images from the validation set of COCO-2014 (Lin et al. 2014) as visual resources. For each image, we rely on crowd-sourcing to manually label paired questions, with one involving a counterfactual presupposition and the other without. Afterward, we leverage GPT-4V to generate three additional counterfactual questions for diversity. Statistically, we have a total of 6,000 questions, with one basic question and four counterfactual questions per image. In order to obtain objective and accurate evaluation metrics, the answers are presented as binary choices, and accuracy serves as the evaluation metric. To eliminate the model’s bias toward preferring a particular choice position, we follow the methodology from previous work (Li et al. 2023a) by adopting the answer ranking strategy for evaluation. Specifically, conditioned on embedded image tokens and the query sentence semantics, we select the choice with highest probability predicted by the MLLM. In addition, we extensively evaluate the prevalent MLLMs on our CFMM dataset and provide further in-depth analyses.
In general, our contributions can be summarized as follows:
  • We curate a novel counterfactual reasoning evaluation benchmark, CFMM, designed to comprehensively assess current MLLMs on counterfactual tasks across diverse aspects, including counting, color, size, shape, direction, and common sense.
  • We provide an objective and accurate evaluation for current state-of-the-art MLLMs, experiment results indicate significant performance degradation for all MLLMs when confronted with counterfactual tasks.
  • We analyze the performance of current state-of-the-art MLLMs on the CFMM dataset and specifically discuss how lightweight techniques such as few-shot learning, ICL, and CoT affect the model’s counterfactual reasoning capabilities, in the hope of providing meaningful support for further research in these areas.
2 Related Works
Multimodal Large Language Models. With the scaling of data size and model size, large language models (LLMs) have shown remarkable performance across various linguistic tasks (Zhang et al. 2022a; Chung et al. 2022; Touvron et al. 2023; Anil et al. 2023). Researchers have delved into extensive explorations to extend the robust generalization capabilities inherent in LLMs into multimodal counterparts (MLLMs). Flamingo (Alayrac et al. 2022) and Open-Flamingo (Awadalla et al. 2023) incorporate two trainable modules, namely the Perceiver Resampler and XAttn-Dense, which enhance the few-shot performance of the model. LLaMA-Adapter (Zhang et al. 2023b; Gao et al. 2023) and Otter (Li et al. 2023c) insert trainable cross-attention layers into LLMs to integrate visual and language information. BLIP2 (Li et al. 2023d) proposes Q-Former to bridge the modality gap between the frozen LLM and vision encoder. InstructBLIP (Dai et al. 2024), MiniGPT-4 (Zhu et al. 2023), VPGTrans (Zhang et al. 2024), mPLUG-OWL (Ye et al. 2023) retain Q-Former while utilizing manually-annotated or GPT4-generated high-quality data for second-stage fine-tuning to enhance the models’ instruction-following abilities. LLaVA (Liu et al. 2024) directly maps visual features to LLM through a lightweight projection layer. Shikra (Chen et al. 2023) and Kosmos-2 (Peng et al. 2023) further enhance MLLMs’ visual grounding abilities such as region description and localization. Besides, Qwen-VL (Bai et al. 2023) increase the input resolution of the visual encoder to 448 × 448, reducing the information loss caused by down-sampling. Monkey (Li et al. 2023e) goes even further by splitting high-resolution images into smaller patches, enabling the model to support resolutions up to 1344×896. In addition, the closed-source business model GPT-4V (Achiam et al. 2023) also demonstrates remarkable image comprehension capabilities. In CFMM, we provide a comprehensive evaluation of these models to thoroughly assess their multimodal counterfactual reasoning capabilities.
Counterfactual Reasoning. “Counterfactual” is an informal term that refers to different perturbations made to factual information. There have been several attempts to assess LLMs’ counterfactual reasoning abilities. CRASS (Frohberg and Binder 2022) proposes the first question answering dataset that requires LLMs to understand the “if” clause consisting of a counterfactual presupposition. IfQA (Yu et al. 2023) explores hypothetical changes of world knowledge in open-domain question answering scenarios. (Li, Yu, and Ettinger 2023) further investigates the impact of cue words and whether the model could infer real-world states from counterfactual conditions. (Wu et al. 2023b) designs counterfactual counterparts for 11 default tasks, and evaluating recent strong LLMs on them. ACQUIRED (Wu et al. 2023a) proposes a counterfactual version of video question answering, where the model is required to infer events that did not actually occur in the video based on the hypothetical language descriptions. VQA-CP (Agrawal et al. 2018) and CF-VQA (Niu et al. 2021) define counterfactual visual question answering as situations where the attribute distributions differ between the training and testing sets. Most related to our work is C-VQA (Zhang et al. 2023a), which leverages ChatGPT to incorporate counterfactual presuppositions into questions sourced from VQAv2 (Goyal et al. 2017). However, constrained by the questions of VQA-v2 and their automated annotation process, its questions conditions lack diversity in question types. Additionally, the answer types are limited to numerical and boolean, which may not fully capture performance in open-world scenarios.
In-Context Learning. LLMs exhibit promising in-context learning (ICL) capabilities (Brown et al. 2020), that is, learning from a few input-label examples from downstream tasks without training. In terms of in-context examples selection, KATE (Liu et al. 2021) retrieves the k-nearest neighbors that are semantically-similar to a test sample. EPR (Min et al. 2022) extends this approach by training a prompt retriever with contrastive loss, which then select demonstrations based on the confidence scores during inference. Auto-CoT (Zhang et al. 2022b) categorizes problems into clusters, employing zero-shot CoT to automatically construct demonstrations for each cluster. Multimodal ICL (M-ICL) is a complex and relatively unexplored field. PICa (Yang et al. 2022) directly translates images into textual descriptions as inputs to GPT-3. MM-React (Yang et al. 2023) additionally incorporates outputs from various visual experts (e.g., dense captioning, object tagging) as inputs to ChatGPT. Science-QA (Lu et al. 2022), MMIC-IT (Li et al. 2023b) and MMICL (Zhao et al. 2023) directly provide demonstrations within the dataset, allowing the model to utilize instruction-tuning to further enhance the capabilities of M-ICL. In this paper, we validated the effectiveness of the in-context learning ability of MLLMs on CFMM to a certain extent.
3 CFMM Benchmark
Figure 2: Demonstration of questions of six types included in our CFMM. For each image, we provide the corresponding basic and counterfactual questions in the first two rows, as well as the answer options in the last row. We include the correct answer of each question in the bracket at the beginning of the question sentence. Note that counterfactual presuppositions may also serve as distractors irrelevant to answer selection as shown in the left case in the last row.
We present the CounterFactual MultiModal reasoning (CFMM) benchmark, a manually curated benchmark to evaluate the counterfactual reasoning capabilities of MLLMs. In this section, we will first elaborate on the definition and taxonomy of counterfactual questions, and then introduce the dataset construction process, as well as evaluation metrics calculation.
3.1 Problem Definition
Counterfactual questions refer to new questions formed by adding imaginary presuppositions to known facts of basic questions. In the NLP field, “facts” are often declarative statements describing background information, “presuppositions” are hypothetical alterations to this information, and “questions” often involve content related to causal inferences. This paradigm extends into the multimodal domain, where “facts” are substituted with multimodal elements like images or videos. For MLLM, answering counterfactual questions essentially requires conditioning on presuppositions and facts to output the answer with the highest probability. We formulate it as a function $f$: $X$$Y$ that maps the input $x\in X$ to the output $y\in Y$:
$f(v,w_{cf},w_{q})=\mathop{\arg\max}\limits_{y^{\prime}}\mathbf{P}(y^{\prime}|v,w_{cf},w_{q}).$
Here, $y^{\prime}$ is the output of MLLM obtained through an appropriate decoder. $v$ and $w_{cf}$, $w_{q}$ represent the image, imaginary presuppositions, and basic question respectively.
Dataset Data Source Question Source Question Types #Questions
Textual Question Answering
CRASS ATOMIC Human - 274
IfQA Wikipedia Human - 3803
Counterfactual Reasoning Psycholinguistic Experiments Automatic 3 400
Reasoning or Reciting NLP Datasets Automatic 11 -
Visual Question Answering
C-VQA-Real VQAv2 Automatic 2 3144
CFMM COCO Human & Auto. 6 6000
Table 1: Statistical comparison with existing counterfactual reasoning benchmarks in both Text-QA and VQA tasks.
3.2 Diverse Aspects of Counterfactual Abilities
Different presuppositions demand disdinct capabilities, thus we progressively challenge MLLMs by setting tasks across 6 aspects:
  • Count: Calculate changes in the quantity of objects when specific items are added to or removed from a group. This challenges the model’s ability to understand numerical operations.
  • Color: Infer how the color of objects changes when they are exchanged or merged. This requires the model to grasp concepts of color mixing and transformation.
  • Shape: Analyze how objects deform or change shape when external forces are applied or when they interact with other objects. This tests the model’s understanding of spatial transformations and physical properties.
  • Size: Imagine changes in the size of objects in space and compare them with other objects. This challenges the model’s ability to reason about spatial relationships and quantitative comparisons of magnitude.
  • Direction: Envision changes in the orientation or position of objects. The model needs to understand changes in spatial location.
  • Common: Change world conditions and ask the model to reason based on external knowledge. This can encompass a wide range of questions that require the model to apply logic and knowledge to new scenarios.
Examples of each question type are illustrated in figure 2. This hierarchical division aims to assess MLLMs’ ability to comprehend and infer entities, attributes, and relationships within images at different levels.
3.3 Dataset Construction
Data Source. Counterfactual question involves hypothetical conditions that entail changes in different aspects, which requires images to contain rich scenarios and potential semantic information for hypothesis. The COCO dataset (Lin et al. 2014) is widely recognized for its extensive image quantity and its focus on object detection, image captioning, and scene understanding. We chose to annotate images from its validation set because of the diverse object categories and rich visual scenes it offers, which capture complex everyday scenarios closely aligned with real-world distributions. This makes COCO particularly well-suited for assessing models’ multimodal counterfactual reasoning capabilities in open-world scenarios. While our images are sourced from COCO, all questions and answers are all newly annotated, ensuring the fairness of our benchmark as the models’ training won’t include these question-answer pairs, even if they had seen the COCO images before.
Manually question annotation. We ask annotators to perform counterfactual question annotations, which follows these steps: Given an image, annotators first select an appropriate question type. Images with no suitable question types will be excluded. Following the selected question type, annotators must initially craft an basic question to assess the models’ basic visual comprehension ability, then modify certain conditions within it to create a counterfactual question. To further enhance the data quality of CFMM, we follow a strict data quality control protocols. Specifically, each question is verified by at least one of the paper’s authors. We primarily examine the following two key points:
  1. (1)
    Information Leakage refers to the situation that answer to a question appearing in the “If” clause, for example, “If I painted this bus blue, what color wouold it be?” In this case, the model can derive answers directly from the text without paying attention to the image content.
  2. (2)
    The ambiguity of the answer refers to the inability to derive a precise answer to a question from the image. For example, it is not appropriate to ask about the number of elephants when only a small part of an elephant’s body is visible. We manually modified or removed all annotations with the above issues.
Automated question expansion. We expanded our dataset based on 1,200 manually annotated question pairs. Leveraging GPT-4’s powerful multimodal understanding and reasoning capabilities, we prompted GPT to creatively generate new content by mimicking our manually annotated data. Specifically, we provided GPT with the original image, a basic question, a counterfactual question, and their respective answers. We then instructed GPT to modify the counterfactual condition while retaining the core question based on the image, thereby generating three new counterfactual questions. After this process, we ultimately obtained 6,000 questions. The prompts used for each category varied slightly, with the detailed prompts provided in the supplementary material. The newly generated questions were required to align with the image information and logical consistency, ensuring that the answers remain as concise and definitive as possible. We also incorporated manual verification to guarantee question quality in this stage.
3.4 Evaluation Strategy
Choice Extraction. Given the challenge of quantitatively evaluating open-ended answers, we structure the answers in the form of common multiple-choice questions with two candidate answers. Considering that the existing instruction-following capabilities of MLLMs are limited. Even if explicitly asked to answer “A” or “B”, the models may still produce some free-form text. Furthermore, we also observe that many MLLMs exhibit a preference for certain options, tending to favor those presented earlier. We adopt the same answer ranking strategy as SEED-Bench-2(Li et al. 2023a) to extract the model’s selection. Specifically, given a question, we compute the generation loss for each option and select the minimum loss as the model’s predicted result. This strategy eliminate the influence of option preferences and instructions, allowing for a more accurate reflection of the model’s actual capabilities.
Evaluation Metrics. We utilize the accuracy and accuracy+ proposed in MME(Fu et al. 2023a), where accuracy here represent correct responses to basic or counterfactual questions, and the accuracy+ indicates that both questions were answered correctly. The random accuracy rates for these two metrics are 50% and 25%, respectively. It can be seen that accuracy+ is a stricter metric, reflecting the extent to which the model fully comprehends the image. Unlike MME, we directly utilize the sum of accuracy and accuracy+ across all questions as our metric. Therefore, the full score of $\text{acc}_{\text{b}}$, $\text{acc}_{\text{c}}$ and $\text{acc}_{\text{+}}$ are all 600.
Model Prompt Count Color Size Shape Direction Common $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ $\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$ Qwen-VL 0-shot 86.78 50.65 42.24 89.14 74.43 67.99 69.23 52.05 34.23 90.09 51.35 42.57 69.62 54.43 33.86 78.79 58.18 41.97 1-shot ICL 86.49 64.44 55.24 87.33 76.47 70.25 69.23 61.92 41.41 89.19 57.43 48.65 71.52 55.70 34.34 78.79 69.70 52.12 1-shot CoT 89.37 62.14 53.02 90.95 76.13 69.80 69.35 60.64 40.77 85.59 55.18 46.62 67.09 57.28 35.60 77.81 67.27 49.85 ChatTruth 0-shot 87.36 70.33 61.49 92.31 76.70 70.93 75.38 55.77 41.53 82.88 58.78 44.59 74.68 49.53 35.60 76.97 66.82 48.03 1-shot ICL 86.21 69.54 60.70 92.31 76.58 70.81 73.84 55.51 41.41 82.88 59.46 45.27 71.52 50.63 37.18 77.58 66.36 47.58 1-shot CoT 87.65 68.61 59.70 90.05 76.36 70.59 74.35 53.72 39.87 81.98 59.23 45.50 70.88 50.00 36.87 73.33 66.67 47.88 LLaVA 0-shot 87.36 50.65 41.88 90.05 71.27 64.03 59.49 53.85 30.64 74.77 69.59 48.65 66.46 56.33 33.23 75.15 64.85 43.94 1-shot ICL 85.64 53.66 45.19 90.05 73.64 66.63 58.98 54.23 30.90 74.77 70.50 50.23 65.19 55.70 33.70 76.97 65.91 45.15 1-shot CoT 81.04 53.66 45.04 87.34 71.95 64.93 56.41 56.15 31.67 76.57 71.17 50.45 63.3 54.59 32.75 73.94 64.70 43.79 InfMLLM 0-shot 91.09 63.94 57.61 93.21 72.40 67.42 66.15 55.38 36.28 78.38 66.44 50.00 74.68 55.06 36.08 82.42 68.03 52.88 1-shot ICL 92.53 61.42 54.89 89.14 72.29 67.31 66.12 55.13 35.90 79.28 65.32 48.87 75.31 54.91 36.71 84.24 67.12 52.42 1-shot CoT 89.08 60.13 53.52 91.85 72.29 67.31 70.77 56.15 36.79 77.48 67.12 50.90 74.05 54.75 36.71 81.21 68.18 53.33 InstructBLIP 0-shot 84.77 67.39 56.61 90.50 72.74 65.72 58.97 56.79 31.28 79.28 55.86 41.89 56.33 53.64 29.59 80.61 61.52 48.79 1-shot ICL 86.78 51.44 42.67 90.5 52.38 46.83 58.66 48.46 28.72 78.48 38.96 28.60 55.7 47.63 26.27 78.39 38.94 28.94 1-shot CoT 80.75 48.20 40.37 85.98 54.64 48.87 59.02 48.85 28.21 81.08 38.06 28.83 50.63 49.53 25.63 78.72 39.24 27.88 BLIP2 0-shot 74.43 53.66 37.72 81.45 69.91 57.13 66.67 65.77 42.95 73.87 52.25 32.21 53.16 52.06 23.26 79.39 61.67 47.58 1-shot ICL 74.71 42.74 28.74 81.9 54.64 45.02 63.59 47.82 28.97 72.97 35.14 20.95 51.26 45.25 22.15 78.18 38.48 29.09 1-shot CoT 72.99 49.93 34.77 81.01 52.94 44.34 66.67 49.62 30.64 71.17 38.29 24.32 56.96 41.61 20.89 81.21 40.61 30.76 Monkey 0-shot 88.79 61.14 53.30 92.76 73.42 68.10 69.74 58.59 39.23 88.29 52.70 42.57 74.22 50.32 34.81 84.33 64.55 52.42 1-shot ICL 89.65 60.06 52.30 92.31 73.30 68.10 69.23 57.44 38.85 88.29 49.55 39.86 75.32 49.21 33.23 84.85 64.55 52.73 1-shot CoT 87.64 59.77 51.87 91.73 73.19 67.87 68.71 57.44 38.97 89.75 51.80 42.12 76.59 49.37 33.54 83.98 65.45 53.64
Table 2: Performance of prevalent MLLMs on six tasks within our proposed CFMM. Here, $\text{acc}_{\text{b}}$ represents the accuracy for correctly answering basic questions, $\text{acc}_{\text{c}}$ denotes the accuracy for correctly answering counterfactual questions, and $\text{acc}_{\text{+}}$ signifies the accuracy for both types of questions. We highlight the best and second best results with red and blue, respectively.
Model Prompt Total score
$\text{acc}_{\text{b}}$ $\text{acc}_{\text{c}}$ $\text{acc}_{\text{+}}$
Qwen-VL 0-shot 483.65 341.10 262.86
1-shot ICL 482.55 385.66 302.01
1-shot CoT 480.16 378.64 295.65
ChatTruth 0-shot 489.58 377.93 302.17
1-shot ICL 484.34 378.08 302.95
1-shot CoT 478.23 374.59 300.41
LLaVA 0-shot 453.27 366.53 262.37
1-shot ICL 451.60 373.64 271.79
1-shot CoT 438.60 372.22 268.63
InfMLLM 0-shot 485.94 381.25 300.27
1-shot ICL 486.62 376.18 296.10
1-shot CoT 484.44 378.61 298.57
InstructBLIP 0-shot 450.46 367.93 273.88
1-shot ICL 448.51 277.80 202.03
1-shot CoT 436.18 278.52 199.79
BLIP2 0-shot 428.97 355.32 240.83
1-shot ICL 422.61 264.08 174.92
1-shot CoT 430.01 272.99 185.72
Monkey 0-shot 498.13 360.72 290.43
1-shot ICL 499.65 354.11 285.07
1-shot CoT 498.40 357.02 288.01
Table 3: The overall performances of prevalent MLLMs on CFMM. $\text{acc}_{\text{b}}$, $\text{acc}_{\text{c}}$ and $\text{acc}_{\text{+}}$ have the same meaning with those in Table 2. The “Total score” indicates the summation of scores achieved on six tasks. We highlight the best and second best results with red and blue, respectively.
3.5 Dataset Analysis
In Table 1, we list all datasets relevant to counterfactual reasoning. It can be observed that existing datasets primarily focus on Text QA, with relatively fewer explorations into VQA. The dataset most similar to ours is C-VQA. However, the types of questions and answers in C-VQA are limited, and we also superior in terms of scale, CFMM can more comprehensively evaluates the counterfactual ability of the model from different aspects. Moreover, we provide basic-counterfactual questions set for each image, which can more intuitively reflect whether the model has truly engaged in counterfactual reasoning. Therefore, the CFMM we propose is a high-quality, comprehensive dataset for evaluating the counterfactual ability of the model.
4 Evaluation Results
In this section, we evaluate the performance of current state-of-the-art open source MLLMs on CFMM. We evaluated a total of 7 open source MLLMs, including Qwen-VL (Bai et al. 2023), ChatTruth, Monkey  (Li et al. 2023e), LLaVA (Liu et al. 2024), InfMLLM (Zhou et al. 2023), InstructBLIP (Dai et al. 2024) and BLIP2 (Li et al. 2023d). During the evaluation, we also tried two training-free prompt-tuning methods, 1-shot ICL and 1-shot CoT (Wei et al. 2022), to explore their impact on the counterfactual ability of MLLMs. The CoT prompts are based on the proven approach outlined in ScienceQA (Lu et al. 2022), utilizing manually annotated CoT processes as in-context learning examples. Our evaluation results exhibit several interesting findings about these MLLMs and hope that these findings will be helpful for the development of the MLLMs community.
4.1 Main Results
Our CFMM has posed formidable challenges to all existing state-of-the-art MLLMs. To date, no model has been able to effectively meet these challenges, and their performance under counterfactual presuppositions has proven unreliable. The evaluation results of MLLMs across different question types and overall scores are presented in Table 2 and Table 3 respectively. It can be seen that all MLLMs perform well on basic questions, achieving an accuracy of around 80%. For “color” questions, InfMLLM even reaches an accuracy of 93.21%. By comparison, a noticeable decline can be observed when evaluating counterfactual questions. All models perform close to the level of random guessing (50%) in the 0-shot scenario. LLaVA exhibited the largest performance drop on counting problems, with an accuracy of only 50.63% on counterfactuals, but an accuracy of 87.36% on basic problems. Overall, Monkey performs best on basic questions, demonstrating the strongest multimodal understanding capability. InfMLLM, due to the introduction of extra VQA datasets while training, possesses stronger query comprehension ability, hence achieving better performance on counterfactual questions.
4.2 Observations
Based on the above experimental results, we summarized four key observations, analyzed their potential causes, and suggest directions for future research.
Observation 1: Eyes can deceive. All MLLMs face significant performance degradation when dealing with counterfactual presuppositions.
All MLLMs exhibited a significant performance drop when confronted with counterfactual questions, with an average decrease of 105.75 points in total score compared to $\text{acc}_{\text{b}}$. Notably, our analysis through $\text{acc}_{\text{+}}$ revealed a slight negative correlation coefficient in the models’ performance on basic questions and counterfactual questions, which indicates that when a model correctly answers a basic question, it is more likely to have committed an error on a counterfactual question. Likewise, the converse holds. This suggests a noteworthy phenomenon: even when a model provides correct answers to counterfactual questions, it may not have grasped the underlying counterfactual presuppositions. Instead, it might have simply relied on the wrong visual cues it obtained and chosen the “answer” it thought corresponding to the basic question. In most cases, visual information is accurate, and models rely on this information while ignoring counterfactual presuppositions, leading to a significant performance drop on $\text{acc}_{\text{c}}$.
Observation 2: The evaluated MLLMs are most sensitive to single-object attributes and worst sensitive to relative relation attributes.
Table2 indicates that MLLMs perform best on basic questions involving single-object attributes such as count and color, but perform worst on questions involving multiple-object attributes related to spatial positional relations, such as size and direction. Existing MLLMs mostly utilize independent visual encoders like CLIP for extracting visual information, yet CLIP and similar encoders are trained for static short descriptions (Fu et al. 2023b; Radford et al. 2021). This makes CLIP and similar visual encoders most sensitive to object recognition and object-level visual features, but less sensitive to spatial positional relations. As can be seen from the results, MLLMs exhibit a significant performance drop in counterfactual questions about counting and color, with $\text{acc}_{\text{c}}$ dropping by over 20 points. When querying the visual information output by an independent visual encoder, object attributes are easiest obtained, but if the MLLM does not understand the counterfactual presuppositions properly, the queried visual information may instead lead to misleading answers.
Figure 3: Qualitative results of “Outputted a smooth but illogical CoT”.
Observation 3: The improvement brought by 1-shot ICL is inconsistent.
Experimental results show that introducing 1-shot ICL into the counterfactual reasoning task significantly improved Qwen-VL’s total score by more than 40 points.
The introduction of 1-shot ICL appears to help MLLM better understand and utilize contextual information, enabling it to more accurately capture key details of counterfactual presuppositions in the reasoning process. However, it is worth noting that while 1-shot ICL has brought some improvements, it is not effective for all MLLMs. For instance, 1-shot ICL has basically no effect on models such as ChatTruth and InfMLLM that are optimized with additional data, and the performance noticeably drops in models that have undergone additional instruction tuning, such as InstructBLIP, or in smaller scale models like BLIP2. Previous studies have shown that the few-shot ICL performance improves as the model’s parameter scale increases (Wei et al. 2023), which may explain why 1-shot ICL did not perform well on the MLLMs we evaluated. Further explorations of the effectiveness of few-shot ICL in counterfactual reasoning on open-source MLLMs are also worth considering.
Observation 4: 1-shot CoT does not help much on MLLMs at 7B-level.
Similar to 1-shot ICL, 1-shot CoT does not significantly enhance the counterfactual reasoning ability of 7B-level MLLMs and may even lead to performance drop in some cases. In Wei et al.’s study (Wei et al. 2022), only models of a certain scale, specifically those exceeding 100B parameters, demonstrated advanced performance when combined with CoT strategies. CoT forces models to engage in reasoning, yet 7B-level models often lack sufficient foundational knowledge and reasoning capabilities, making it unlikely for CoT to generate correct reasoning steps and potentially leading to issues like AI hallucinations. For instance, Figure 3 shows an output example produced by Qwen-VL 7B when using CoT. It can be seen that the model outputs fluent but illogical CoT, resulting in incorrect results. Ben et al. (Prystawski, Li, and Goodman 2024) also mentioned in their research that when using CoT, the model’s training data should be task-relevant and interconnected. CoT demonstrates excellent performance when the training data of large models exhibit local clusters of variables. However, visual counterfactual reasoning data is absent in the training datasets of MLLMs. Therefore, it is necessary to explore how CoT can be effectively implemented in smaller-scale MLLMs to reduce the cost of practical applications, enabling MLLMs to be more widely utilized.
5 Conclusion
In this work, we introduce CFMM, a novel benchmark designed for evaluating the counterfactual reasoning capabilities of multimodal large language models (MLLMs). CFMM consists of 1,200 image and 6,000 question with accurate human annotation or verification covering 6 assessment dimensions. We conduct a comprehensive evaluation of 7 well-known open source MLLMs, analyze and compare their performance to provide insights for future research. We plan to launch and maintain a leaderboard to provide the community with a reliable platform to evaluate MLLMs’ counterfactual abilities.
Appendix A Dataset Statistics
CFMM comprises 1200 images, each accompanied by one basic question and four counterfactual question. Fig. 4(a) provides a detailed breakdown of the number of questions within each type. Note that despite our requirement for each annotator to provide balanced annotations for every question type, there still exists some imbalance among them. Specifically, “count” are easier to formulate as counterfactuals, thus constituting a larger proportion of questions. In contrary, “if” conditions in “shape” often results in information leakage, rendering them more challenge to annotate. The distributions of the length of questions are shown in Fig. 4(c) and 4(d). Counterfactual questions necessitate the addition of “if” conditions to basic questions, resulting in an average length of 17.97 words—much longer than that of basic questions, which is 9.51 words. Fig. 4(b) display the word cloud of answers. For answers, we aim for simplicity and clarity, hence their average length is only 1.28 words.
(a) Question types and corresponding number.
(b) Word cloud of answers.
(c) Distribution of word lengths in basic questions.
(d) Distribution of word lengths in counterfactual questions.
Figure 4: Detailed statistics of the proposed CFMM. The quantitative distribution of six types of questions is shown in (a), and the frequency of words in answers is shown in (b). We also list the distribution of the length of basic questions and counterfactual questions in (c) and (d), respectively.
Appendix B Details of automated question expansion
We used the version gpt-4-turbo-2024-04-09 for generation, which was trained on data up to December 2023. To ensure that GPT generates new counterfactual questions that meet our requirements, we needed to tailor its focus and considerations for each category. Therefore, we employed customized system prompts for different categories.
For categories count, color and common, we used the following prompt:
{mdframed}
You are now a question extension assistant for a counterfactual reasoning problem. Given an image and a basic VQA question with its answer, rewrite the counterfactual question provided and generate three new counterfactual questions and answers. The new counterfactuals should alter the original counterfactual condition but retain the core question. Let your imagination guide the generation while maintaining logical possibilities related to the original context, ensure that the question should be as concise as possible and the answer to the new counterfactual questions should be as certain as possible. The format of input and output is as follows, output the new questions using the following JSON structure:
Input: [Image here] “category”: “[Category type]”, “basic question”: “[Original basic question]”, “basic answer”: “[Original basic answer]”, “counterfactual question”: “[Original counterfactual question]”, “counterfactual answer”: “[Original counterfactual answer]”
Output: “category”: “[Category type]”, “basic question”: “[Original basic question]”, “basic answer”: “[Original basic answer]”, “counterfactual question”: “[Original counterfactual question]”, “counterfactual answer”: “[Original counterfactual answer]”, “counterfactual question1”: “[New counterfactual question 1]”, “counterfactual answer1”: “[Answer to new counterfactual question 1]”, “counterfactual question2”: “[New counterfactual question 2]”, “counterfactual answer2”: “[Answer to new counterfactual question 2]”, “counterfactual question3”: “[New counterfactual question 3]”, “counterfactual answer3”: “[Answer to new counterfactual question 3]”
For categories size and direction, we used the following prompt:
{mdframed}
You are now a question extension assistant for a counterfactual reasoning problem. Given an image and a basic VQA question with its answer, rewrite the counterfactual question provided and generate three new counterfactual questions and answers. The new counterfactuals should alter the original counterfactual condition but retain the core question. Let your imagination guide the generation while maintaining logical possibilities related to the original context, ensure that the question should be as concise as possible. Please generate new counterfactual questions that have clear answers based on the information in the provided image. The format of input and output is as follows, output the new questions using the following JSON structure:
Input: [Image here] “category”: “[Category type]”, “basic question”: “[Original basic question]”, “basic answer”: “[Original basic answer]”, “counterfactual question”: “[Original counterfactual question]”, “counterfactual answer”: “[Original counterfactual answer]”
Output: “category”: “[Category type]”, “basic question”: “[Original basic question]”, “basic answer”: “[Original basic answer]”, “counterfactual question”: “[Original counterfactual question]”, “counterfactual answer”: “[Original counterfactual answer]”, “counterfactual question1”: “[New counterfactual question 1]”, “counterfactual answer1”: “[Answer to new counterfactual question 1]”, “counterfactual question2”: “[New counterfactual question 2]”, “counterfactual answer2”: “[Answer to new counterfactual question 2]”, “counterfactual question3”: “[New counterfactual question 3]”, “counterfactual answer3”: “[Answer to new counterfactual question 3]”
For category common, we used the following prompt:
{mdframed}
You are now a question extension assistant for a counterfactual reasoning problem. Given an image and a basic VQA question with its answer, rewrite the counterfactual question provided and generate three new counterfactual questions and answers. The new counterfactuals should alter the original counterfactual condition but retain the core question. Let your imagination guide the generation while maintaining logical possibilities related to the original context, ensure that the question should be as concise as possible. Please generate new counterfactual questions with clear answers based on the information in the provided image, along with your world knowledge and common sense. The format of input and output is as follows, output the new questions using the following JSON structure:
Input: [Image here] “category”: “[Category type]”, “basic question”: “[Original basic question]”, “basic answer”: “[Original basic answer]”, “counterfactual question”: “[Original counterfactual question]”, “counterfactual answer”: “[Original counterfactual answer]”
Output: “category”: “[Category type]”, “basic question”: “[Original basic question]”, “basic answer”: “[Original basic answer]”, “counterfactual question”: “[Original counterfactual question]”, “counterfactual answer”: “[Original counterfactual answer]”, “counterfactual question1”: “[New counterfactual question 1]”, “counterfactual answer1”: “[Answer to new counterfactual question 1]”, “counterfactual question2”: “[New counterfactual question 2]”, “counterfactual answer2”: “[Answer to new counterfactual question 2]”, “counterfactual question3”: “[New counterfactual question 3]”, “counterfactual answer3”: “[Answer to new counterfactual question 3]”